### v1

# Adaptive Coordinate-Wise Step Sizes for Quasi-Newton Methods: A Learning-to-Optimize Approach

Abstract
Tuning effective step sizes is crucial for the stability and efficiency of optimization algorithms. While adaptive coordinate-wise step sizes tuning methods have been explored in first-order methods, second-order methods still lack efficient techniques. Current approaches, including hypergradient descent and cutting plane methods, offer limited improvements or encounter difficulties in second-order contexts. To address these challenges, we introduce a novel Learning-to-Optimize (L2O) model within the Broyden-Fletcher-Goldfarb-Shanno (BFGS) framework, which leverages neural networks to predict optimal coordinate-wise step sizes. Our model integrates a theoretical foundation that establishes conditions for the stability and convergence of these step sizes. Extensive experiments demonstrate that our approach achieves substantial improvements over traditional backtracking line search and hypergradient descent-based methods, offering up to 7$\times$ faster and stable performance across diverse optimization tasks.
1 Introduction
Step size is an essential hyperparameter in optimization algorithms. It determines the rate at which the optimization variables are updated, and greatly influences the convergence speed and stability of the optimization process. In first-order gradient-based optimization, how to choose an appropriate step size is well studied: The step size is typically adjusted adaptively using past gradient information such as in AdaGrad [15], RMSProp [16], and Adam [17] for large-scale deep learning. Despite differences in details, these adaptive methods assign a single scalar step size to all optimization variables (e.g. model parameters in ML) which works well in practice.
Step size in second-order methods received much less attention thus far. Second-order methods leverage the curvature information to adjust both the search direction and step size, offering faster convergence in number of iterations for high-dimensional optimization landscapes, at the cost of high computational complexity in calculating the Hessian (or its approximation). A natural and common approach for step size selection here is line search, which iteratively adjusts the step size along the descent direction until certain conditions, like the Armijo condition, are met [3].
In contrast, we study the more general coordinate-wise step sizes in this work, which allow for individual variables to have different step sizes. Coordinate-wise step sizes are beneficial since different optimization variables may have different sensitivities to the step size; scalar step size is obviously a special case of coordinate-wise step sizes. They have also been shown to improve convergence in first-order methods [1, 18, 15].
In this work, we explore the impact of coordinate-wise step sizes in the context of second-order methods, which remains largely unexplored to our knowledge. We choose the Broyden-Fletcher-Goldfarb-Shanno (BFGS) method [8], one of the most widely used second-order optimization methods, as the backbone method. BFGS belongs to the quasi-Newton family of methods that iteratively update an approximation of the Hessian matrix using gradient information to reduce the complexity.
We start our study by demonstrating that existing solutions to tune coordinate-wise step sizes in first-order methods do not work well in second-order contexts. The first such approach is hypergradient descent [21, 22], which iteratively tunes step sizes using their gradients at each BFGS step. We show empirically that it provides only marginal gains after the initial few steps of BFGS. Moreover, cutting-plane techniques, which expand backtracking line search into multiple dimensions, iteratively refine step sizes within feasible sets narrowed down by hypergradient-based incisions [18]. This method essentially offers an approximation of the Hessian in a first-order framework, thus complicating its direct application to second-order methods, in which Hessian approximation is handled by BFGS update, and the step sizes are adjusted to improve the Hessian approximation. Further, the intricate curvature within the Hessian presents additional challenges in plane cutting.
Therefore, we explore the Learning to Optimize (L2O) paradigm [2] in this work. L2O replaces handcrafted rules with data-driven models that can adaptively learn efficient strategies, tailoring optimization processes to specific problem structures [2, 20]. L2O has shown promising results in first-order optimization by leveraging neural networks to predict optimal step sizes dynamically based on the current optimization state [19, 26].
The application of L2O in second-order methods presents challenges. Whereas in first-order approaches, the step size primarily regulates the update magnitude, in second-order methods, it also affects the precision of Hessian approximations, particularly in the initial optimization stages. This dual role adds complexities in tuning the step size. Furthermore, the unconstrained exploration inherent in conventional L2O makes convergence and stability harder to achieve within second-order L2O frameworks.
To address these challenges, we provide a theoretical analysis of coordinate-wise step sizes within the BFGS framework. We establish specific convergence criteria and derive conditions that effective step sizes must satisfy. This analysis identifies key elements of the optimization state that influence step size selection and provides bounded conditions on step sizes that ensure convergence. Additionally, we derive that the coordinate-wise step sizes should asymptotically converge to an identity matrix to achieve super-linear convergence rates near the optimum.
Building on this theoretical foundation, we propose a customized L2O model for second-order optimization. Our model takes as input variables, gradients, and second-order search directions. To meet the derived convergence conditions, we bound the output of the L2O model. Unlike typical first-order L2O methods that use fixed unrolling lengths, we train our model with a loss function defined by the expected objective value in the subsequent iteration plus a regularization term, ensuring efficient adaptation of step sizes.
We summarize our key contributions as follows:
  1. 1.
    We are the first to investigate coordinate-wise step size adaptation in the context of second-order optimization methods, specifically the BFGS algorithm.
  2. 2.
    We derive conditions that effective coordinate-wise step sizes should satisfy, ensuring convergence and stability based on desired optimization properties.
  3. 3.
    We propose a new L2O method to generate coordinate-wise step sizes for the BFGS algorithm, integrating both theoretical principles and adaptive learning to guide the optimization process.
  4. 4.
[…]
2 Preliminaries
In this chapter, we introduce the basics of second-order optimization methods, with a focus on BFGS. We show how step size selection critically affects both the convergence and the quality of Hessian approximations. Then we establish the key assumptions that will support our analysis of coordinate-wise step sizes in BFGS optimization.
2.1 Second-Order Methods
Second-order optimization methods, such as Newton’s method [4], utilize both gradient and curvature information to find the minimum of an objective function. While first-order methods typically achieve a sub-linear convergence rate [6], second-order methods generally exhibit a faster, quadratic convergence rate [28]. This significant difference makes second-order methods particularly advantageous in high-dimensional optimization problems, allowing for more efficient progress towards the optimum [5].
In Newton’s method, the objective function is locally approximated by a quadratic function around the current parameter vector $x_{k}$:
$g(y)\approx f(x_{k})+\nabla f(x_{K})^{T}(y-x_{k})+\frac{\alpha_{k}}{2}(y-x_{k})^{T}H_{k}(y-x_{k}),$
where $H_{k}$ is the Hessian matrix and $\alpha_{k}$ is the damped parameter. By minimizing the quadratic approximation, the update rule for Newton’s method becomes [28]:
$x_{k+1}=x_{k}-\alpha_{k}H_{k}^{-1}\nabla f(x_{k}).$
Computing the Hessian is quite expensive and often infeasible for large-scale problems [23]. Instead, quasi-Newton method was proposed to approximate the Hessian to be more affordable and scalable [14, 9]. Generally, quasi-Newton methods maintain an approximation of the Hessian matrix $B_{k}\approx H_{k}$ at each iteration, updating it with a rank one or rank two term based on the gradient differences between two consecutive iterations [11, 8]. During this process, the Hessian approximation is restricted to follow the secant equation [28]:
$B_{k+1}s_{k}=y_{k},$ (1)
where $s_{k}=x_{k+1}-x_{k}$ and $y_{k}=\nabla f(x_{k+1})-\nabla f(x_{k})$. In the most common BFGS method, the Hessian approximation $B_{k}$ is updated at each iteration using the formula:
$B_{k+1}=B_{k}-\frac{B_{k}s_{k}s_{k}^{T}B_{k}}{s_{k}^{T}B_{k}s_{k}}+\frac{y_{k}y_{k}^{T}}{y_{k}^{T}s_{k}},$
and the update rule becomes:
$x_{k+1}=x_{k}-\alpha_{k}B_{k}^{-1}\nabla f(x_{k}).$
Although the enrollment of curvature information can greatly assist the optimization process, it also makes the algorithm more sensitive to the step size selection [27]. The step size influences the update of the Hessian approximation, and an inappropriately large step can lead to violations of the curvature condition $y_{k}^{T}s_{k}>0$, potentially resulting in an indefinite Hessian approximation [28].
The quality of the Hessian approximation is also heavily dependent on the step size. If $\alpha_{k}$ is too small, the changes in gradients $y_{k}$ and parameters $s_{k}$ may be seriously influenced by numerical errors, leading to poor Hessian updates and slower convergence [13]. In quasi-Newton methods, the step size must balance between exploiting the current curvature information (encoded in $B_{k}$) and allowing for sufficient exploration of the parameter space. This balance is more delicate than in first-order methods due to the adaptive nature of the search direction.
2.2 Assumptions
Our objective is to minimize the convex objective function $f(x)$ over $x\in\mathbb{R}^{n}$:
$\min_{x\in\mathbb{R}^{n}}f(x).$
For our analysis, we make the following assumptions on the objective function and the Hessian approximations, which are widely regarded as reasonable and commonly used in optimization research [26, 19, 28]:
Assumption 1.
The objective function $f$ is $L$-smooth, meaning there exists a constant $L$ such that:
$\|\nabla^{2}f(x)-\nabla^{2}f(y)\|\leq L\|x-y\|.$
Assumption 2.
The gradient $\nabla f(x)$ is differentiable in an open, convex set $D$ in $\mathbb{R}^{n}$, and $\nabla^{2}f(x)$ is continuous at $x^{*}$ with $\nabla^{2}f(x^{*})$ being nonsingular.
Assumption 3.
The Hessian approximation generated by BFGS method is positive definite and has a uniformly bounded condition number. Specifically, there exists a constant $M$ such that:
$\lambda_{\min}(B_{k})\geq\frac{1}{M}\lambda_{\max}(B_{k}),$
where $\lambda_{\min}(B_{k})$ and $\lambda_{\max}(B_{k})$ are the smallest and largest eigenvalues of $B_{k}$, respectively. By this assumption we assume the Hessian approximation is not too ill-conditioned.
Assumption 4.
The norm of update direction $B_{k}^{-1}\nabla f(x_{k})$ is upper bounded by a constant $R$:
$\|B_{k}^{-1}\nabla f(x_{k})\|\leq R.$
This assumption is reasonable, as quasi-Newton method maintain boundedness of $B_{k}^{-1}$ through stable Hessian approximations [9], and gradients $\nabla f(x_{k})$ typically diminish near optimal points, ensuring the update direction remains controlled.
3 Coordinate-Wise Step Sizes for BFGS
In this section, we first analyze the theoretical advantages of coordinate-wise step sizes and then explore hypergradient descent as a practical method for coordinate-wise step sizes tuning. However, the limited improvements achieved through hypergradient descent reveal the challenges of finding effective coordinate-wise step sizes, prompting us to consider alternative approaches. We resort to L2O method that can directly learn the step sizes from data derived from similar optimization problems. Building on this perspective, we establish sufficient conditions for effective coordinate-wise step sizes that ensure convergence and descent properties, thus laying a solid foundation for learning-based approaches that can predict optimal step sizes efficiently during optimization.
3.1 Gain of Coordinate-Wise Step Sizes
To illustrate the potential benefits of coordinate-wise step sizes in the BFGS method, let us consider the theoretical implications of relaxing the constraint that step size should be a scalar. Assume we have identified an optimal scalar step size, denoted by $\alpha_{k}^{*}$, for the $k$-th iteration. Since the restriction of a convex function to a line remains convex [7], this optimal step size $\alpha_{k}^{*}$ satisfies:
$\frac{d}{d\alpha_{k}}f(x_{k+1})\bigg{|}_{\alpha_{k}=\alpha_{k}^{*}}$ (2)
$=$ $\frac{d}{d\alpha_{k}^{*}}f(x_{k}-\alpha_{k}^{*}B_{k}^{-1}\nabla f(x_{k}))$
$=$ $-\nabla f(x_{k}-\alpha_{k}^{*}B_{k}^{-1}\nabla f(x_{k}))^{\top}B_{k}^{-1}\nabla f(x_{k})$
$=$ $0$
When constrained to a scalar form, $\alpha_{k}^{*}$ guarantees optimality along the single search direction $B_{k}^{-1}\nabla f(x_{k})$. However, if we allow the step size to be a diagonal matrix $P_{k}$ rather than a scalar, the optimality condition of $\alpha_{k}^{*}$ may no longer hold. By extending to a coordinate-wise approach, we aim to further minimize the objective function by adjusting each coordinate independently, which can potentially achieve a lower function value than with $\alpha_{k}^{*}$ alone.
To explore this, let $P_{k}=\alpha_{k}^{*}I$, and consider the partial derivative of $f(x_{k+1})$ with respect to $P_{k}$ at this point:
$\frac{\partial}{\partial P_{k}}f(x_{k}-P_{k}B_{k}^{-1}\nabla f(x_{k}))\bigg{|}_{P_{k}=\alpha_{k}^{*}I}$ $=$ (3)
$\text{diag}(-\nabla f(x_{k}-\alpha_{k}^{*}B_{k}^{-1}\nabla f(x_{k}))\odot$ $B_{k}^{-1}\nabla f(x_{k})),$
where $\odot$ denotes the Hadamard (element-wise) product. Since $B_{k}^{-1}\nabla f(x_{k})\neq 0$ , the derivative in equation 3 equals zero only if $\nabla f(x_{k}-\alpha_{k}^{*}B_{k}^{-1}\nabla f(x_{k}))=0$. Since the optimum does not generally lie on the direction of $B_{k}^{-1}\nabla f(x_{k})$, the dot product being zero in equation 2 does not imply that the Hadamard product is also zero in equation 3. This observation suggests that even we know the optimal scalar step size $\alpha^{*}$, we can still find coordinate-wise step sizes that could achieve a more effective descent.
To determine suitable coordinate-wise step sizes, we employ hypergradient descent. Defining $g(p)=f(x_{k}-p\odot B_{k}^{-1}\nabla f(x_{k}))$, where $p$ is the diagonal of $P$,we can analyze the smoothness of $g(p)$ as follows:
$\quad\|\nabla g(p_{1})-\nabla g(p_{2})\|$
$=\|(\nabla f(x_{k}-p_{1}\odot B_{k}^{-1}\nabla f(x_{k}))$
$\quad\quad-\nabla f(x_{k}-p_{2}\odot B_{k}^{-1}\nabla f(x_{k}))\|$
$\leq L\|(p_{1}-p_{2})\odot B_{k}^{-1}\nabla f(x_{k})\|$
$\leq L\|B_{k}^{-1}\nabla f(x_{k})\|\|p_{1}-p_{2}\|$
$\leq LR\|p_{1}-p_{2}\|,$
where $L$ is the Lipschitz constant of $\nabla f$ and $R$ is from assumption 4. This shows that $g(p)$ is $LR$-smooth. Based on this property, we can set the coordinate-wise step sizes $P_{k}$ as:
$P_{k}=\alpha_{k}^{*}I-\frac{1}{LR}v_{k}B_{k}^{-1}\nabla f(x_{k}),$
where $v_{k}=\operatorname{diag}(\nabla f(x_{k}-\alpha_{k}^{*}B_{k}^{-1}\nabla f(x_{k})))$. This coordinate-wise step size $P_{k}$ is theoretically guaranteed to perform better than the scalar step size $\alpha_{k}^{*}$:
$\begin{split}&f(x_{k}-P_{k}B_{k}^{-1}\nabla f(x_{k}))\leq f(x_{k}-\alpha_{k}^{*}B_{k}^{-1}\nabla f(x_{k}))\\ &\quad-\frac{1}{2LM}|\nabla f(x_{k}-\alpha_{k}^{*}B_{k}^{-1}\nabla f(x_{k}))\odot B_{k}^{-1}\nabla f(x_{k})|^{2}.\end{split}$
This demonstrates that coordinate-wise step sizes in the BFGS method can yield a more substantial decrease in the objective function than a scalar step size.
3.2 Numerical Analysis of Coordinate-Wise Step Size: A Hypergradient Descent Method
HGD (i) 1 5 10 20
BFGS (k)
1 7.52938 6.32887 5.22274 4.32551
2 1.97834 1.95869 1.93509 1.89111
3 0.88499 0.88143 0.87703 0.86839
4 0.44807 0.44746 0.44670 0.44519
5 0.25669 0.25658 0.25644 0.25617
6 0.14995 0.14992 0.14988 0.14979
7 0.08901 0.08900 0.08899 0.08896
8 0.05425 0.05425 0.05424 0.05423
9 0.03373 0.03373 0.03372 0.03372
10 0.02107 0.02107 0.02107 0.02107
Table 1: Objective value of the least square problem with hypergradient descent (HGD) on $P_{k}$ for different BFGS iterations.
Coordinate-wise step sizes operate in a higher-dimensional search space compared to traditional line search methods, which are confined to a single dimension. Given $B_{k}^{-1}\nabla f(x_{k})$, any vector in the space can be achieved as $P_{k}B_{k}^{-1}\nabla f(x_{k})$ through an appropriate choice of $P_{k}$. There exists an optimal $P_{k}^{*}$ such that $P_{k}^{*}B_{k}^{-1}\nabla f(x_{k})=x-x^{*}$, allowing the optimization to reach the solution in a single iteration. However, finding this optimal $P_{k}^{*}$ is as hard as finding $x^{*}$ itself. Therefore, we seek a computationally efficient approach to find the coordinate-wise step sizes $P_{k}$ that can enhance convergence.
Building upon section 3.1, we investigate hypergradient descent on coordinate-wise step size matrix $P_{k}$. The BFGS update rule with coordinate-wise step size takes the form:
$x_{k+1}=x_{k}-P_{k}B_{k}^{-1}\nabla f(x_{k}).$ (4)
We initializes $P_{k}^{0}$ as the identity matrix $I$ before each BFGS update [10]. We then perform hypergradient descent on $P_{k}$ using the gradient of $f(x_{k+1})$ with respect to $P_{k}^{i}$ to obtain $P_{k}^{i+1}$. After $T$ iterations, we employ $P_{k}^{T}$ in the BFGS update rule 4. The hypergradient descent on $P_{k}$ can be formulated as:
$P_{k}^{i+1}=P_{k}^{i}-\eta\frac{\partial f(x_{k}-P_{k}^{i}B_{k}^{-1}\nabla f(x_{k}))}{\partial P_{k}^{i}},$
where $\eta$ is the learning rate for the gradient descent on $P_{k}$.
We conduct experiments on the least squares problem to assess the effectiveness of hypergradient descent applied to $P_{k}$. Each BFGS iteration includes 20 steps of hypergradient descent, after which the most recent $P_{k}$ identified by hypergradient descent is used in BFGS udate. Table 1 presents the experimental results, where each row shows the objective value within one BFGS iteration across different hypergradient descent steps. The results demonstrate that while hypergradient descent shows some improvement over standard BFGS, the benefits become increasingly marginal as iterations progress. This implies that finding an effective $P_{k}$ is inherently challenging. From an alternating optimization perspective, convergence to $x^{*}$ can be achieved through updates to either $P_{k}^{i}$ or $x_{k}$. Hypergradient descent on $P_{k}^{i}$, despite being computationally lightweight as a first-order method, offers limited performance gains. In contrast, BFGS updates, while computationally more demanding due to their need to maintain accurate Hessian approximations, provide a more substantial optimization progress.
It is reasonable to allocate the majority of computational resources to the BFGS process itself due to its fast convergence. For the selection of $P_{k}$, we then seek a method that can provide meaningful improvements without incurring significant computational costs. While no direct formula exists for obtaining the optimal $P_{k}$, and hypergradient descent, despite its theoretical guarantees, requires multiple iterations to yield satisfactory results, we envision a more efficient approach. This leads us to consider a question: Can we leverage the patterns in optimization trajectories to generate effective step sizes directly? In many optimization scenarios, similar patterns of gradients and Hessian approximations may warrant similar step size adjustments. If these patterns could be learned from data, we might be able to bypass the iterative computation entirely. This insight motivates us to explore the L2O paradigm, which use neural networks to predict $P_{k}$ directly.
L2O has shown strong potential in capturing complex patterns and relationships, making it suitable for tasks like predicting step sizes based on optimization state features. By leveraging a neural network, we could potentially map the current optimization state—possibly using gradient information and Hessian approximations—to coordinate-wise step sizes directly. This approach would allow immediate predictions of effective step sizes without iterative refinement.
Our L2O model, introduced in Section 4, is built with these objectives in mind. But before developing our learning-based approach, we establish theoretical foundations by identifying key requirements for effective coordinate-wise step sizes in BFGS optimization. These requirements characterize the essential properties that guide the design of our L2O method.
3.3 Conditions for Good Coordinate-Wise Step Sizes
Effective coordinate-wise step size must satisfy certain theoretical requirements to ensure that each iterative update is convergent. In this section, we establish the foundational conditions required for these coordinate-wise step sizes to achieve convergence properties. These conditions provide a systematic framework for the development of any adaptive coordinate-wise step size tuning methods that maintain the stability and efficiency of BFGS method.
For good coordinate-wise step sizes, we propose the following requirements:
  1. 1.
    The generated sequence ${x_{k}}$ converges to one of the local minimizers of $f$.
  2. 2.
    Each update moves towards the minimizer.
  3. 3.
    The method achieves superlinear convergence.
The first requirement extends the concept of Fixed Point Encoding in optimization, serving as a valuable guideline for constructing effective optimization algorithms [25]. The second requirement emphasizes directional accuracy, ensuring each update consistently progresses toward the minimizer. Finally, the third requirement preserves the superlinear convergence rate characteristic of the BFGS method [28].
We now present three theorems that provide sufficient conditions for coordinate-wise step sizes to satisfy the proposed requirements.
Theorem 1.
Let $\{x_{k}\}$ be the sequence generated by equation 4. If the coordinate-wise step size $P_{k}$ satisfies
$\|P_{k}\|$ $\leq\frac{\alpha}{L\|B_{k}^{-1}\|},$
$\|P_{k}\|$ $\geq\beta\frac{\nabla f(x_{k})^{\top}B_{k}^{-1}\nabla f(x_{k})}{\|B_{k}^{-1}\nabla f(x_{k})\|^{2}},$
for certain $20<\alpha<$ 2 and $\beta>0$, where $L$ is the Lipschitz constant of gradients and $B_{k}$ is the approximate Hessian generated by BFGS, then the sequence of gradients converges to zero: $\lim_{k\to\infty}\|\nabla f(x_{k})\|=0$.
Proof.
See Appendix 7.1. ∎
Remark.
Theorem 1 establishes robust sufficient conditions for gradient convergence while maintaining substantial implementation flexibility. The theorem’s bounds on $P_{k}$ are particularly accommodating: the lower bound can be made arbitrarily close to zero through appropriate selection of $\beta$, while setting $\alpha$ near $2$ allows the upper bound to approach $2/(L|B_{k}^{-1}|)$, which remains strictly less than 2 if $|B_{k}^{-1}|>L$. This is true since $B_{k}$ is the average of Hessians from $x_{k-1}$ to $x_{k}$, as evidenced by equation 1.
Theorem 1 suggests a pragmatic simplification: constraining the elements of the coordinate-wise step size to the interval [0,2] should be sufficient for practical implementations. Moreover, theorem 1 indicates that $P_{k}$ should be computed as a function of both the gradient $\nabla f(x_{k})$ and Hessian approximation $B_{k}$, as evidenced by the presence of both gradient and Hessian information in bounds. Notably, these results extend beyond convex optimization, requiring only L-smoothness of the objective function rather than convexity.
Theorem 2.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a twice continuously differentiable convex function that has Lipschitz continuous gradient with $L>0$. Let $x^{*}$ denote the unique minimizer of $f$. Suppose that $\{B_{k}\}$ is a sequence of approximate Hessians such that they are uniformly lower bounded:
$\gamma I\preceq B_{k},$
for certain constant $\gamma>0$. Let $\{P_{k}\}$ be a sequence of diagonal matrices with entries $p_{k,i}$ satisfying:
$0<p_{k,i}\leq\frac{2\gamma}{L},$
Define the iterative sequence $\{x_{k}\}$ by equation 4. Then, the sequence $\{x_{k}\}$ satisfies:
$\|x_{k+1}-x^{*}\|\leq\|x_{k}-x^{*}\|.$
Proof.
See Appendix 7.2. ∎
Remark.
Since $B_{k}$ captures the average Hessian behavior between consecutive points $x_{k-1}$ and $x_{k}$, its eigenvalues lie within the bounds of $\nabla^{2}f(x)$, yielding $\gamma\leq L$. This relationship reveals that the seemingly restrictive upper bound $\frac{2\gamma}{L}$ for $p_{k,i}$ simplifies to 2. This aligns with Theorem 1, as both theorems suggest the coordinate-wise step sizes should lie within the interval [0,2]. However, theorem2 makes an additional assumption of convexity, which enables a stronger guarantee, i.e., each iteration strictly decreases the distance to the optimum.
The theorem can be reduced to classical optimization methods in certain scenarios. For instance, setting $B_{k}=I$ and $P_{k}=\alpha I$ with $\alpha\leq 2/L$ yields the standard gradient descent method with constant step size.
Moreover, the proof of theorem 2 indicates that optimal $P_{k}$ values should minimize the spectral radius of $T_{k}=I-P_{k}B_{k}^{-1}H_{k}$. As the algorithm progresses ($k\to\infty$), $B_{k}$ approaches $H_{k}$, suggesting that $P_{k}$ should converge to the identity matrix. This convergence behavior is formally established in the subsequent theorem.
Theorem 3.
Suppose $\nabla f(x)$ is differential in the open, convex set $D$ in $\mathbb{R}^{n}$, and $\nabla^{2}f(x)$ is continuous at $x^{*}$ and $\nabla^{2}f(x^{*})$ is nonsingular. $\{B_{k}\}$ is generated by BFGS and $\{x_{k}\}$ is generated by the update rule. If $\{x_{k}\}$ converges to $x^{*}$. Then $\{x_{k}\}$ converges superlinearly to $x^{*}$ if $\{P_{k}\}$ converges to the identical matrix.
Proof.
See Appendix 7.3. ∎
Theorem 3 provides crucial insight into the asymptotic behavior of coordinate-wise step sizes. When the iterates are far from the optimum, coordinate-wise step sizes can accelerate convergence by adapting to the local geometry of the objective function. However, as the algorithm approaches the optimum, the BFGS method naturally provides increasingly accurate Hessian approximations. At this stage, additional coordinate-wise scaling becomes unnecessary and could potentially interfere with the superlinear convergence properties of BFGS. It suggests that adaptive schemes for $P_{k}$ should be designed to gradually reduce their influence as the optimization progresses, eventually allowing the natural BFGS updates to dominate near the optimum.
4 L2O Model
Building on the theoretical foundations established in the previous section, we now present our L2O model for coordinate-wise step size selection in BFGS optimization. Our method leverages neural networks to efficiently predict step sizes that satisfy the key requirements while adapting to the local geometry of the optimization landscape. This section details the architecture, training procedure, and design choices that enable effective step size prediction across diverse optimization problems.
We propose an L2O method using an LSTM (Long Short-Term Memory) network to predict coordinate-wise step sizes [19]. The architecture is structured as follows:
$h_{k},o_{k}$ $=\text{LSTM}(x_{k},\nabla f(x_{k}),B_{k}^{-1}\nabla f(x_{k}),h_{k-1},\phi_{\text{LSTM}}),$
$p_{k}$ $=\text{MLP}(o_{k},\phi_{\text{MLP}}),$
$P_{k}$ $=\text{diag}(2\sigma(p_{k})),$
where, $h_{k}$ is the LSTM hidden state, initialized randomly for the first iteration, and $o_{k}$ is the embedding output from the LSTM network. The parameters of the LSTM and MLP (Multi-Layer Perceptron) networks are denoted by $\phi_{\text{LSTM}}$ and $\phi_{\text{MLP}}$, respectively. According to section 3.3, the values of $p_{k,i}$ should lie within the interval [0, 2]. Therefore, the activation function $\sigma$, which represents the sigmoid function, is employed to ensure that the predicted step sizes are within a suitable range, enabling bounded and adaptive adjustments for each coordinate.
To enhance scalability and parameter efficiency, we employ a coordinate-wise LSTM approach, where the same network is shared across all input coordinates, as suggested in [2, 20]. This design allows the L2O method to adapt to problems of varying dimensionality without an increase in the number of parameters, making it highly efficient for large-scale optimization tasks.
The L2O model is trained on datasets of diverse optimization problems, allowing it to learn common structures and behaviors across different problem instances. The training process comprises two nested optimization loops that interact as follows:
Inner Optimization Loop
In the inner loop, we utilize the current L2O model to optimize the given objective function. For each training instance, which is an optimization problem, we perform totally $K$ iterations. At each iteration $k$, the L2O model predicts the step size $P_{k}$, which is then used to update the current point. This process allows the L2O model to adaptively adjust the step size for each coordinate based on the current optimization state, effectively utilizing the curvature information encoded in $B_{k}^{-1}$.
Outer Optimization Loop
Immediately after each iteration of the inner loop, we update the parameters of the neural networks ($\phi_{\text{LSTM}}$ and $\phi_{\text{MLP}}$) using backpropagation. The update is based on the objective function value achieved at $x_{k+1}$. Unlike other L2O approaches in first-order context, which often employ a fixed unroll length and update the model after multiple iterations [2, 26, 20], our method updates the model parameters more frequently. This frequent update strategy allows the model to capture the immediate impact of its predictions on the optimization trajectory.
The loss function used for training draws inspiration from the concept of exact line search, a technique in optimization that determines the optimal step size along a given search direction to minimize the objective function precisely. Specifically, we aim to minimize the objective function value in the next iteration with a regularization term:
$\min_{\phi_{\text{LSTM}},\phi_{\text{MLP}}}\mathbb{E}_{f\sim\mathcal{F}}[f(x_{k+1})]+\lambda\|P_{k}-I\|_{F}$
where $\mathcal{F}$ represents the distribution of optimization problems used for training and $\lambda$ is the regularization parameter. The regularization term ensures that as we approach the optimum, the coordinate-wise step sizes converge toward an identical matrix, aligning with the insight from Theorem 3.
In BFGS method, the step size can be viewed as a correction to the Hessian approximation. In early optimization stages, the objective value primarily drives the loss function, and the Hessian approximation may lack precision. Thus, an adaptive coordinate-wise step size is necessary to enhance the accuracy of the Hessian approximation based on the current state. However, as the optimization nears convergence, the Hessian approximation becomes more accurate, shifting the influence on the loss function to the regularization term. At this point, the coordinate-wise step sizes converge to the identity matrix, as further corrections to the Hessian approximation are no longer required.
Figure 1: Comparison on least square.
Figure 2: Comparison on logistic regression with different regularization parameter.
Figure 3: Comparison on log-sum-exp functions with different regularization parameter.
Short Term Dependencies
The rationale behind our frequent update strategy lies in the complex dynamics of the BFGS-based L2O method. The predicted coordinate-wise step size has a dual effect: it influences the subsequent iterate $x_{k+1}$ directly and impacts the update of the BFGS Hessian approximation $B_{k+1}$. This creates a highly intricate feedback loop, where the step size affects not only the optimization trajectory but also the quality of the Hessian approximation, which in turn affects future predictions.
Deferring the neural network update to occur after multiple iterations, as is common in most first-order L2O methods, could result in a significant delay between action and feedback. This delay might obscure the nuanced effects of individual step size choices, impairing the network’s ability to learn effective strategies for optimization. Frequent updates, on the other hand, provide the network with immediate feedback, thereby enhancing its learning capacity and responsiveness.
Long-Term Dependencies
Despite the frequent updates, our model retains the ability to capture long-term dependencies in the optimization process. The use of LSTM networks ensures that the historical information is effectively encoded within the hidden state $h_{k}$. This enables the model to leverage patterns and dependencies across the optimization trajectory, balancing short-term adaptability with long-term strategic planning.
The combination of coordinate-wise step sizes, frequent model updates, and LSTM-based memory allows our L2O approach to effectively navigate the complex optimization landscapes typical in machine learning applications, achieving both fast convergence and robustness against local irregularities in the objective function’s geometry.
5 Experiments
Our experiments are conducted using Python 3.9 and PyTorch 1.12 on an Ubuntu 18.04 system, utilizing an Intel Xeon Gold 5320 CPU and two NVIDIA RTX 3090 GPUs. We employ the Adam optimizer as our meta-optimizer to train our L2O model on a dataset comprising 32,000 optimization problems with randomly sampled parameters. For evaluation, we generate a separate test dataset of 1,024 optimization problems. Our L2O method is benchmarked against two baselines: BFGS with backtracking line search, and BFGS with hypergradient descent. We implement BFGS with backtracking line search by initially setting the step size to 1 and iteratively scaling it by a factor until the Armijo condition is met, ensuring adequate reduction in the objective function. For BFGS with hypergradient descent, we refine the coorindate-wise step sizes by conducting 20 iterations of hypergradient descent within each BFGS iteration. We experiment with various parameter settings for all methods and select the best-performing configurations.
Optimization Objectives
We evaluate our L2O method on three distinct optimization problems: least square, logistic regression, and log-sum-exp function.
For least square, we use the objective function:
$\min_{x}f(x)=\frac{1}{2}\|Ax-b\|^{2},$
where $A\in\mathbb{R}^{250\times 500}$ and $b\in\mathbb{R}^{500}$ are randomly generated using a Gaussian distribution. Following the setting in [19], we set the sparsity of $A$ to 0.1 by making 90% of its element to zero.
The objective function for logistic regression is:
$\min_{x}f(x)=\frac{1}{m}\sum_{i=1}^{m}$ $[b_{i}\log(h(a_{i}^{T}x))$
$+(1-b_{i})\log(1-h(a_{i}^{T}x))]+\rho\|x\|_{2}^{2},$
where $m=500$, $\{(a_{i},b_{i})\in\mathbb{R}\times\{0,1\}\}_{i=1}^{m}$ are randomly generated, $h(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function.
For the log-sum-exp function, we use:
$\min_{x}f(x)=\log\left(\sum_{i=1}^{m}e^{a_{i}^{T}x-b_{i}}\right)+\rho\|x\|^{2}_{2},$
where $m=500$, $\{(a_{i},b_{i})\in\mathbb{R}^{250}\times\mathbb{R}\}_{i=1}^{m}$. We followe the dataset generation process described in [24]. We first generate auxiliary random vectors $\{\hat{a_{i}}\}_{i=1}^{m}$ by sampling uniformly from the interval $[0,1]$. We then generate $\{b_{i}\}_{i=1}^{m}$ from the standard normal distribution. Using these, we define an auxiliary function $\hat{f}(x)=\log\left(\sum_{i=1}^{m}e^{\hat{a}_{i}^{T}x-b_{i}}\right)$. Finally, we set $a_{i}=\hat{a}_{i}-\nabla\hat{f}(0)$, ensuring that the optimal solution of $f(x)$ is 0.
Comparison with Baselines
Figure 2 illustrates the results for the least squares problems, where the optimization target is zero. We terminate the process when the gradient norm drops below $10^{-10}$ As shown, hypergradient descent offers negligible improvement, while our proposed approach achieves a reduction of approximately 20 iterations, equivalent to a 25% improvement in convergence speed. The linear-like trajectory of our method further validates its superlinear convergence.
In figure 2, we present the results for logistic regression with various regularization parameters. All three methods converge rapidly. Although hypergradient descent attains a lower objective value than line search during the process, both methods exhibit similar convergence iteration counts. In contrast, our proposed method accelerates convergence during initial stages and consistently achieves lower objective values throughout.
Finally, figure 3 displays the results on log-sum-exp problems. Our method reaches the optimum in around 10 iterations, while the other methods require 22 to 70 iterations, depending on $rho$. For smaller regularization parameters $\rho$, the problem’s ill-conditioning intensifies as the objective’s Hessian approaches zero. While baseline methods deteriorate significantly under lower $\rho$, our method demonstrates robustness to ill-conditioning, consistently achieving faster convergence than the other two approaches.
6 Conclusions
This work investigates the use of coordinate-wise step sizes in BFGS method. We examine the advantages and challenges of identifying coordinate-wise step sizes by theoretical and numerical analysis. We rigorously derive conditions to enhance convergence properties. Building on these findings, we develop an L2O model that predicts optimal coordinate-wise step sizes. Experimental results demonstrate that our proposed approach outperforms baseline methods in terms of both convergence speed and stability.
Supplementary Material
7 Proofs
7.1 Proof of Theorem 1
Proof.
Consider a quadratic function
$Q(x)=f(y)+\nabla f(y)^{\top}(x-y)+\frac{\alpha}{2}(x-y)^{\top}B_{k}P_{k}^{-1}(x-y).$
Basically, we use this quadratic function as a local approximation of $f(x)$ around $y$, and we think the minimum of this quadratic function is a better solution than $y$. This can work only if $Q(x)$ is an overestimation of $f(x)$. Indeed, we can show that:
$\quad Q(x)$
$=f(y)+\nabla f(y)^{\top}(x-y)+\frac{\alpha}{2}(x-y)^{\top}B_{k}P_{k}^{-1}(x-y)$
$\geq f(y)+\nabla f(y)^{\top}(x-y)+\frac{\alpha}{2}\frac{1}{\|B_{k}^{-1}\|\|P_{k}\|}\|x-y\|^{2}$
$\geq f(y)+\nabla f(y)^{\top}(x-y)+\frac{L}{2}\|x-y\|^{2}$
$\geq f(x),$
where the second inequality uses the condition $\|P_{k}\|\leq\frac{\alpha}{L\|B_{k}^{-1}\|}$ and the third uses the assumption of $L$-smoothness.
Plugging $x=y-P_{k}B_{k}^{-1}\nabla f(y)$ into the inequality, we get the Armijo condition:
$f(y)-(1-\frac{\alpha}{2})\nabla f(y)^{\top}P_{k}B_{k}^{-1}\nabla f(y)\geq f(y-P_{k}B_{k}^{-1}\nabla f(y)).$
Let $x_{k}=y$ and $x_{k+1}=y-P_{k}B_{k}^{-1}\nabla f(y)$, we have
$f(x_{k+1})$ $\leq f(x_{k})-(1-\frac{\alpha}{2})\nabla f(x_{k})^{\top}P_{k}B_{k}^{-1}\nabla f(x_{k})$
$\leq f(x_{k})-(1-\frac{\alpha}{2})\beta\frac{(\nabla f(x_{k})^{\top}B_{k}^{-1}\nabla f(x_{k}))^{2}}{\|B_{k}\nabla f(x_{k})\|^{2}}$
$=f(x_{k})-(\beta-\frac{\alpha\beta}{2})\cos^{2}\theta_{k}\|\nabla f(x_{k})\|^{2},$
where $\theta_{k}$ is the angle between $\nabla f(x_{k})$ and $B_{k}^{-1}\nabla f(x_{k})$. In the second inequality, we use the condition that $\|P_{k}\|\geq\beta\frac{\nabla f(x_{k})^{T}B_{k}^{-1}\nabla f(x_{k})}{\|B_{k}^{-1}\nabla f(x_{k})\|^{2}}$.
Following the proof of Theorem 3.2 in [28], summing over all iterations, we have:
$f(x_{k+1})\leq f(x_{0})-(\beta-\frac{\alpha\beta}{2})\sum_{i=0}^{k}\cos^{2}\theta_{i}\|\nabla f(x_{i})\|^{2}.$
Note that $f(x_{0})-f(x_{k+1})$ is lower-bounded by 00 and upper-bounded by $f(x_{0})$. Hence when $k$ approaches infinity, we have:
$\sum_{i=0}^{k}\cos^{2}\theta_{i}\|\nabla f(x_{i})\|^{2}<\infty,$
which implies
$\cos^{2}\theta_{k}\|\nabla f(x_{k})\|^{2}\to 0.$
Since $\|B_{k}\|\|B_{k}^{-1}\|<M$,
$\cos\theta_{k}$ $=\frac{\nabla f(x_{k})^{\top}B_{k}^{-1}\nabla f(x_{k})}{\|\nabla f(x_{k})\|\|B_{k}^{-1}\nabla f(x_{k})\|}$
$\geq\frac{\lambda_{B_{k}^{-1}}^{\min}\|\nabla f(x_{k})\|^{2}}{\lambda_{B_{k}^{-1}}^{\max}\|\nabla f(x_{k})\|^{2}}$
$=\frac{1/\|B_{k}\|}{\|B_{k}^{-1}\|}$
$>\frac{1}{M}.$
Then we have
$\lim_{k\to\infty}\|\nabla f(x_{k})\|=0.$
7.2 Proof of Theorem 2
Proof.
Let $e_{k}=x_{k}-x^{*}$ denote the difference between the current iterate and the minimizer. We can write the update rule as:
$e_{k+1}=e_{k}-P_{k}B_{k}^{-1}\nabla f(x_{k}).$
Using the mean value theorem for vector-valued functions, We can write the gradient as:
$\nabla f(x_{k})$ $=\nabla f(x_{k})-\nabla f(x^{*})$
[…]
$=H_{k}e_{k}$
where:
[…]
Since $f$ is $L$-smooth , we have:
$\nabla^{2}f(x)\preceq LI.$
Noted that $H_{k}$ is nothing else but the average of the Hessian matrix along the line segment between $x^{*}$ and $x_{k}$. Then we have:
$H_{k}\preceq LI.$
Substitute $\nabla f(x_{k})=H_{k}e_{k}$ into the error update, we have:
$e_{k+1}$ $=e_{k}-P_{k}B_{k}^{-1}H_{k}e_{k}$
$=(I-P_{k}B_{k}^{-1}H_{k})e_{k}.$
Consider the matrix $T_{k}=I-P_{k}B_{k}^{-1}H_{k}$, we will analyze the spectral radius of it. The upper bound of the eigenvalue of $T_{k}$ is 1, since $P_{k}$ is a diagonal matrix with positive entries $p_{k,i}$, and $B_{k}^{-1}$ and $H_{k}$ are positive definite matrices. The lower bound of the eigenvalue of $T_{k}$ is:
$\lambda_{min}(T_{k})$ $=1-\lambda_{max}(P_{k})\lambda_{max}(B_{k}^{-1})\lambda_{max}(H_{k})$
$\geq 1-\frac{2\gamma_{1}}{L}\frac{L}{\gamma_{1}}=-1$
Hence the spectral radius of $T_{k}$ is less than 1. Then we have:
$\|e_{k+1}\|\leq\|T_{k}\|\|e_{k}\|\leq\|e_{k}\|.$
7.3 Proof of Theorem 3
Proof.
We first show that if
$\lim_{k\to\infty}\frac{\|[B_{k}P_{k}^{-1}-\nabla^{2}f(x^{*})](x_{k+1}-x_{k})\|}{\|x_{k+1}-x_{k}\|}=0,$ (5)
then $\{x_{k}\}$ converges to $x^{*}$ superlinearly.
We can write the numerator as
$[B_{k}P_{k}^{-1}-\nabla^{2}f(x^{*})](x_{k+1}-x_{k})$
$=-\nabla f(x_{k})-\nabla^{2}f(x^{*})(x_{k+1}-x_{k})$
$=\nabla f(x_{k+1})-\nabla f(x_{k})$
$\quad-\nabla^{2}f(x^{*})(x_{k+1}-x_{k})-\nabla f(x_{k+1})$
$=\nabla^{2}f((1-\xi)x_{k+1}+\xi x_{k})(x_{k+1}-x_{k})$
$\quad-\nabla^{2}f(x^{*})(x_{k+1}-x_{k})-\nabla f(x_{k+1}),$
where the last equation uses the mean value theorem and $\xi\in(0,1)$.
Since $\{x_{k}\}$ converges to $x^{*}$, we have:
$\lim_{k\to\infty}\frac{\nabla^{2}f((1-\xi)x_{k+1}+\xi x_{k})(x_{k+1}-x_{k}))}{x_{k+1}-x_{k}}$
$\quad-\frac{-\nabla^{2}f(x^{*})(x_{k+1}-x_{k})}{{}_{k+1}-x_{k}}$
$=\lim_{k\to\infty}\nabla^{2}f((1-\xi)x_{k+1}+\xi x_{k})-\nabla^{2}f(x^{*})=0.$
Hence,
$\lim_{k\to\infty}\frac{\|\nabla f(x_{k+1})\|}{\|x_{k+1}-x_{k}\|}=0.$
Since $\nabla^{2}f(x^{*})$ is nonsingular, there exists a $\beta>0$ such that
$\|\nabla f(x_{k+1})\|$ $=\|\nabla f(x_{k+1})-\nabla f(x^{*})\|$
$=\|\nabla^{2}f(x_{\xi})(x_{k+1}-x^{*})\|$
$\geq\beta\|x_{k+1}-x^{*}\|.$
Therefore,
$\frac{\nabla f(x_{k+1})}{\|x_{k+1}-x_{k}\|}$ $\geq\frac{\beta\|x_{k+1}-x^{*}\|}{\|x_{k+1}-x_{k}\|}$
$\geq\frac{\beta\|x_{k+1}-x^{*}\|}{\|x_{k+1}-x^{*}\|+\|x_{k}-x^{*}\|}$
$=\beta\frac{\frac{\|x_{k+1}-x^{*}\|}{\|x_{k}-x^{*}\|}}{1+\frac{\|x_{k+1}-x^{*}\|}{\|x_{k}-x^{*}\|}}.$
Hence, $\lim_{k\to\infty}\frac{\|x_{k+1}-x^{*}\|}{\|x_{k}-x^{*}\|}=0$, which indicates superlinear convergence.
Then we show that equation 5 is indeed true. Theorem 3.4 in [12] shows that the $\{B_{k}\}$ generated by BFGS satisfies
$\lim_{k\to\infty}\frac{\|B_{k}-\nabla^{2}f(x^{*})\|}{\|x_{k+1}-x_{k}\|}=0.$
If $P_{k}$ converges to the identity matrix,
$\quad~{}\lim_{k\to\infty}\frac{\|[B_{k}P_{k}^{-1}-\nabla^{2}f(x^{*})](x_{k+1}-x_{k})\|}{\|x_{k+1}-x_{k}\|}$
$=\lim_{k\to\infty}\frac{\|[B_{k}-\nabla^{2}f(x^{*})](x_{k+1}-x_{k})\|}{\|x_{k+1}-x_{k}\|}$
$\leq\lim_{k\to\infty}\frac{\|B_{k}-\nabla^{2}f(x^{*})\|\|(x_{k+1}-x_{k})\|}{\|x_{k+1}-x_{k}\|}$
$=0.$
And this limitation is obviously non-negative, hence it is zero. Then we can conclude that $\{x_{k}\}$ converges to $x^{*}$ superlinearly. ∎